# Unraveling the Impact of Heterophilic Structures on Graph Positive-Unlabeled Learning

### v1

@@ 5.3 Ablation Study
GPL $\bf{75.1{\scriptstyle\pm 0.9}}$ $\bf{83.3{\scriptstyle\pm 0.8}}$ $\bf{41.9{\scriptstyle\pm 1.3}}$ $\bf{37.2{\scriptstyle\pm 1.4}}$
- w/o LPL ${70.0\downarrow 7.3\%}$ ${79.6\downarrow 4.6\%}$ ${13.9\downarrow 66.8\%}$ ${17.5\downarrow 53.0\%}$
- w/o Bilevel ${71.5\downarrow 4.8\%}$ ${80.1\downarrow 3.8\%}$ ${32.4\downarrow 22.7\%}$ ${28.9\downarrow 22.3\%}$
- w/o Selected ${71.1\downarrow 5.3\%}$ ${80.2\downarrow 3.7\%}$ ${34.6\downarrow 17.4\%}$ ${32.4\downarrow 12.9\%}$
@@ Acknowledgements
TLL is partially supported by the following Australian Research Council projects: FT220100318, DP220102121, LP220100527, LP220200949, and IC190100031. JCY was supported by the National Key R&D Program of China (No. 2022ZD0160703), 111 plan (No. BP0719010) and National Natural Science Foundation of China (No. 62306178). The authors would give special thanks to Suqin Yuan and Muyang Li for helpful discussions and comments. The authors thank the reviewers and the meta-reviewer for their helpful and constructive comments on this work.
@@ F.3 Performance with different GNN architectures
GPL+GCN ${81.9{\scriptstyle\pm 0.5}}$ ${80.1{\scriptstyle\pm 0.4}}$ ${75.1{\scriptstyle\pm 0.9}}$ $\bf{83.3{\scriptstyle\pm 0.8}}$ ${41.9{\scriptstyle\pm 1.3}}$ ${37.2{\scriptstyle\pm 1.4}}$ ${40.7{\scriptstyle\pm 5.8}}$ ${51.0{\scriptstyle\pm 4.2}}$ $\bf{50.8{\scriptstyle\pm 2.7}}$ ${52.3{\scriptstyle\pm 3.2}}$
GPL+GAT $\bf{82.1{\scriptstyle\pm 0.5}}$ $\bf{80.8{\scriptstyle\pm 0.7}}$ ${75.7{\scriptstyle\pm 0.2}}$ ${82.9{\scriptstyle\pm 0.3}}$ ${40.3{\scriptstyle\pm 1.1}}$ ${35.7{\scriptstyle\pm 0.9}}$ $\bf{40.9{\scriptstyle\pm 4.6}}$ ${50.8{\scriptstyle\pm 3.5}}$ ${49.8{\scriptstyle\pm 1.8}}$ $\bf{52.6{\scriptstyle\pm 2.1}}$
GPL+ARMA ${81.2\scriptstyle\pm 1.1}$ ${79.9{\scriptstyle\pm 0.9}}$ ${75.1{\scriptstyle\pm 0.5}}$ ${82.7{\scriptstyle\pm 0.7}}$ $\bf{42.0{\scriptstyle\pm 0.8}}$ ${36.9{\scriptstyle\pm 1.2}}$ ${39.9{\scriptstyle\pm 4.5}}$ ${51.0{\scriptstyle\pm 4.2}}$ ${49.9{\scriptstyle\pm 2.3}}$ ${52.5{\scriptstyle\pm 2.9}}$
GPL+APPNP ${81.5{\scriptstyle\pm 0.6}}$ ${80.3{\scriptstyle\pm 0.5}}$ $\bf{75.9{\scriptstyle\pm 0.7}}$ ${83.1{\scriptstyle\pm 0.6}}$ ${41.7{\scriptstyle\pm 1.8}}$ $\bf{37.8{\scriptstyle\pm 1.1}}$ ${40.1{\scriptstyle\pm 4.9}}$ $\bf{51.5{\scriptstyle\pm 3.8}}$ ${50.1{\scriptstyle\pm 2.9}}$ ${51.9{\scriptstyle\pm 2.9}}$

### v2

@@ 5.3 Ablation Study
GPL $\bf{74.1{\scriptstyle\pm 0.9}}$ $\bf{82.3{\scriptstyle\pm 0.8}}$ $\bf{37.9{\scriptstyle\pm 1.3}}$ $\bf{36.2{\scriptstyle\pm 1.4}}$
- w/o LPL ${70.0\downarrow 5.5\%}$ ${79.6\downarrow 3.3\%}$ ${13.9\downarrow 63.3\%}$ ${17.5\downarrow 51.6\%}$
- w/o Bilevel ${71.5\downarrow 3.5\%}$ ${80.1\downarrow 2.7\%}$ ${32.4\downarrow 14.5\%}$ ${28.9\downarrow 20.2\%}$
- w/o Selected ${71.1\downarrow 5.3\%}$ ${80.2\downarrow 3.7\%}$ ${34.6\downarrow 8.7\%}$ ${32.4\downarrow 10.5\%}$
@@ Acknowledgements
TLL is partially supported by the following Australian Research Council projects: FT220100318, DP220102121, LP220100527, LP220200949, and IC190100031. JCY is supported by the National Key R&D Program of China (No. 2022ZD0160703), 111 plan (No. BP0719010) and National Natural Science Foundation of China (No. 62306178). BH is supported by the NSFC General Program No. 62376235 and Guangdong Basic and Applied Basic Research Foundation Nos. 2022A1515011652 and 2024A1515012399. The authors would give special thanks to Suqin Yuan and Muyang Li for helpful discussions and comments. The authors thank the reviewers and the meta-reviewer for their helpful and constructive comments on this work.
@@ F.3 Performance with different GNN architectures
GPL+GCN ${81.9{\scriptstyle\pm 0.5}}$ ${79.1{\scriptstyle\pm 0.4}}$ ${74.1{\scriptstyle\pm 0.9}}$ $\bf{82.3{\scriptstyle\pm 0.8}}$ ${37.9{\scriptstyle\pm 1.3}}$ ${36.2{\scriptstyle\pm 1.4}}$ ${37.7{\scriptstyle\pm 5.8}}$ ${48.0{\scriptstyle\pm 4.2}}$ $\bf{45.8{\scriptstyle\pm 2.7}}$ ${46.3{\scriptstyle\pm 3.2}}$
GPL+GAT $\bf{82.1{\scriptstyle\pm 0.5}}$ $\bf{79.8{\scriptstyle\pm 0.7}}$ ${74.7{\scriptstyle\pm 0.2}}$ ${81.9{\scriptstyle\pm 0.3}}$ ${36.3{\scriptstyle\pm 1.1}}$ ${35.7{\scriptstyle\pm 0.9}}$ $\bf{37.9{\scriptstyle\pm 4.6}}$ ${48.8{\scriptstyle\pm 3.5}}$ ${44.8{\scriptstyle\pm 1.8}}$ $\bf{46.6{\scriptstyle\pm 2.1}}$
GPL+ARMA ${81.2\scriptstyle\pm 1.1}$ ${78.9{\scriptstyle\pm 0.9}}$ ${74.1{\scriptstyle\pm 0.5}}$ ${81.7{\scriptstyle\pm 0.7}}$ $\bf{38.0{\scriptstyle\pm 0.8}}$ ${35.9{\scriptstyle\pm 1.2}}$ ${36.9{\scriptstyle\pm 4.5}}$ ${47.0{\scriptstyle\pm 2.2}}$ ${44.9{\scriptstyle\pm 2.3}}$ ${46.5{\scriptstyle\pm 2.9}}$
GPL+APPNP ${81.5{\scriptstyle\pm 0.6}}$ ${79.3{\scriptstyle\pm 0.5}}$ $\bf{74.9{\scriptstyle\pm 0.7}}$ ${82.1{\scriptstyle\pm 0.6}}$ ${37.7{\scriptstyle\pm 1.8}}$ $\bf{36.8{\scriptstyle\pm 1.1}}$ ${37.1{\scriptstyle\pm 4.9}}$ $\bf{49.5{\scriptstyle\pm 3.8}}$ ${45.1{\scriptstyle\pm 2.9}}$ ${45.9{\scriptstyle\pm 2.6}}$